# NeuMaDiff: Neural Material Synthesis via Hyperdiffusion

### v1

@@ A.2 Diffusion Model
During inference, the network enables sampling via an iterative process [47], leveraging the factorization of the learned distribution as
@@
for $p({\bm{\mathbf{x}}}_{T}):=\mathcal{N}(0,I)$. For conditional sampling, we employ classifier-free guidance (CFG) [24] (for more details, please refer to Sec. B.3).
A.3 Attention Mechanism
The attention mechanism [52] allows models to focus on specific parts of input data, dynamically assigning different levels of “attention” or importance to different elements. Attention mechanisms enable models to learn which parts of the input sequence are most relevant to predicting each output token. This process improves performance by allowing models to prioritize relevant information and ignore irrelevant details, especially in long sequences.
In modern applications, attention modules are widely used in natural language processing (NLP), computer vision, and beyond. The transformer architecture [52], which relies heavily on the self-attention mechanism, has become foundational in NLP models, including BERT [11] and GPT [33]. Attention allows these models to capture complex dependencies between words in a sentence, regardless of their distance from each other, leading to significant advancements in tasks like language translation, sentiment analysis, and image processing.
The attention mechanism computes a weighted combination of values based on the relevance of each value to a given query. In the context of self-attention (or scaled dot-product attention) in transformer models [52], this process involves three main components: queries (${\bm{\mathbf{q}}}$), keys (${\bm{\mathbf{k}}}$), and values (${\bm{\mathbf{v}}}$). In summary, the attention mechanism dynamically focuses on relevant parts of the input by computing similarity scores between queries and keys, normalizing these scores, and using them to weight the values. This allows models to capture dependencies across elements in a sequence, making attention a powerful tool for handling long-range dependencies in data.
@@ A.3 Attention Mechanism
where $W_{\bm{\mathbf{q}}}$, $W_{\bm{\mathbf{k}}}$, and $W_{\bm{\mathbf{v}}}$ are learnable weight matrices. These projections represent the queries, keys, and values, respectively. The attention score between each query and key is then computed as the dot product ${\bm{\mathbf{q}}}\cdot{\bm{\mathbf{t}}}$. This results in a matrix of scores that represents the similarity between each element in the sequence. To stabilize gradients and prevent large values in the dot-product computation, the scores are scaled by the square root of the dimensionality of the queries/keys $\sqrt{d_{\bm{\mathbf{k}}}}$, where $d_{\bm{\mathbf{k}}}$ is the dimension of the key vectors. The scaled scores are $\frac{{\bm{\mathbf{q}}}\cdot{\bm{\mathbf{t}}}}{\sqrt{d_{k}}}$.
@@
A.4 Variational Autoencoder
We adopt variational autoencoders (VAEs) [28] as one of the baseline models in Sec. 4.3. VAEs are probabilistic generative models designed to capture the underlying probability distribution of a given dataset $\mathcal{D}$.
@@ A.4 Variational Autoencoder
$\mathcal{L}_{\text{ELBO}}({\bm{\mathbf{\psi}}},{\bm{\mathbf{\zeta}}};{\bm{\mathbf{x}}}):=$ $\mathbb{E}{q_{{\bm{\mathbf{\psi}}}}({\bm{\mathbf{z}}}|{\bm{\mathbf{x}}})}\left[\log p_{{\bm{\mathbf{\zeta}}}}({\bm{\mathbf{x}}}|{\bm{\mathbf{z}}})\right]$ (21)
@@
where $\mathcal{D}_{\text{KL}}$ represents the Kullback-Leibler divergence between the approximate posterior and the prior distribution [9]. This probabilistic framework enables VAEs to effectively approximate the true data distribution, facilitating robust generative modeling of complex data.
A.5 K-Means Clustering
We adopt K-means clustering in constrained synthesis, where unconditional materials are filtered out by statistical information rather than a neural network. K-means clustering is a commonly used unsupervised method to partition $n$ data samples ${\bm{\mathbf{X}}}\in\mathbb{R}^{n\times p}$ into $k$ clusters $\mathcal{S}=\{S_{1},...,S_{k}\}$, minimizing the distance between each sample and its center
@@ A.5 K-Means Clustering
B Model Details
B.1 Neural Field
The sum $D_{\text{NF}}$ of dimensionalities of the weights in our neural field $f_{r}^{\bm{\mathbf{\xi}}}$ is 675.
B.2 Transformer Backbone in Hyperdiffusion
@@
Fig. 11 illustrates the transformer architecture.
@@
B.3 Conditional Sampling with Classifier-Free Guidance
We employ classifier-free guidance (CFG) [24] for conditional sampling on the hyperdiffusion. We present the algorithm in Alg. 1. Notice that we require $\omega>-1$ for conditional generation. Otherwise, the model downgrades to unconditional synthesis when $\omega=-1$.
@@ B.3 Conditional Sampling with Classifier-Free Guidance
${\bm{\mathbf{x}}}_{T}\sim\mathcal{N}(0,I_{D_{\text{NF}}})$triangleright sample ${\bm{\mathbf{x}}}_{T}$ from prior
for $t=T$ to $1$ do
@@
B.4 PCA-Based Baselines
For PCA-based baselines introduced in the unconditional synthesis in Sec. 4.3, the dimensionalities of the reduced spaces for PCA-A and PCA-N are 300 and 100, respectively.
B.5 VAE-Based Baselines
In this section, we introduce the model and training hyperparameters for VAE-based baselines introduced in the unconditional synthesis in Sec. 4.3.
For VAE-A, to reduce the input complexity, we downsample it from $90\times 90\times 180$ to $45\times 45\times 90$ and upsample the synthetic results using nearest neighbor algorithm. The hidden dimension is 256 and the latent space dimension is 300. The model is trained with batch size 16, learning rate $5\times 10^{-3}$ until convergence.
For VAE-N, the VAE architecture contains MLP-based encoder and decoder, each with 4 layers. The input dimension is $D_{\text{NF}}$, and the hidden and the latent space dimensions are both 100. The likelihood of the reconstructed data is measured by the mean squared error (MSE). The model is trained with batch size 32, learning rate $10^{-3}$ until convergence.
C Experiment Details
C.1 Training Details
In the neural field fitting, we set the batch size to be 512, the number of epochs to be 100, and the learning rate to be $5\times 10^{-3}$.
When training the hyperdiffusion model, we set the total timestep $T$ to be 100, the batch size to be 512, the number of epochs to be 700 for unconditional synthesis and an additional 200 for conditional synthesis, and the learning rate to be adaptive from $5\times 10^{-4}$ to $5\times 10^{-6}$.
C.2 Full Material Type List
We include the full list of 48 material types used in the type-conditional synthesis in Sec. 4.4: acrylic, alum-bronze, alumina-oxide, aluminium, aventurnine, brass, wood, chrome-steel, chrome, colonial-maple, color-changing-paint, delrin, diffuse-ball, fabric, felt, fruitwood, grease-covered-steel, hematite, ipswich-pine, jasper, latex, marble, metallic-paint, natural, neoprene-rubber, nickel, nylon, obsidian, oxidized-steel, paint, phenolic, pickled-oak, plastic, polyethylene, polyurethane-foam, pvc, rubber, silicon-nitrade, soft-plastic, special-walnut, specular-fabric, specular-phenolic, specular-plastic, stainless-steel, steel, teflon, tungsten-carbide, two-layer
C.3 Full Set of Constrained Synthesis Rules
@@
    |$\lvert\{({\bm{\mathbf{\omega}}}_{o},{\bm{\mathbf{\omega}}}_{i})\in\mathbb{R}^{6}\mid\left\|f_{r}({\bm{\mathbf{\omega}}}_{o},{\bm{\mathbf{\omega}}}_{i})\right\|_{\infty}>f_{\text{diffuse}}\}\rvert<e,$ (24)
[…]
@@
  • Specular materials: Through our statistical analysis, we identify two specular thresholds $f_{\text{specular}}^{(1)}:=100$ and $f_{\text{specular}}^{(2)}:=600$. The materials can be classified as low-, mid-, or high-specular if the maximum reflectance value fall in the range $[f_{\text{diffuse}},f_{\text{specular}}^{(1)}),[f_{\text{specular}}^{(1)},f_{\text{specular}}^{(2)})$, and $[f_{\text{specular}}^{(2)},\infty)$, respectively.
@@
  • Mirror-like materials: the specular lobe in the polar diagram is narrow. In order to quantify this, we observe that under the Rusinkiewicz reparametrization (Sec. 3.2), the evaluation at $f_{r}(\theta_{{\bm{\mathbf{h}}}}=0,\theta_{{\bm{\mathbf{d}}}}=0,\varphi_{{\bm{\mathbf{d}}}}=0)$ is the peak of the lobe. Incrementing $\theta_{\bm{\mathbf{h}}}$ decreases the BRDF. The width $w$ of the lobe can then be defined as the value of $\theta_{\bm{\mathbf{h}}}$ for which the reflectance value drops to half of the the peak $\frac{f_{r}(0,0,0)}{2}$. Empirically, we found that if the width $w<w_{\text{mirror}}:=0.349$, the material exhibits mirror-like behavior.
D Metric Details
D.1 Rendering-Based Metrics
@@
SSIM [54] is a perception-based metric that measures the perceptual similarity of the two images. The computation of SSIM is based on three comparison measurements between the two images: luminance ($l$), contrast ($c$), and structure ($s$) defined as
$l(I_{1},I_{2}):=\frac{2\mu_{x}\mu_{y}+c_{1}}{\mu_{x}^{2}+\mu_{y}^{2}+c_{1}};$ (26)
@@
are the variables to stabilize the division with weak denominator where $k_{1}$ and $k_{2}$ are coefficients defaulted to 0.01 and 0.03, respectively, and $L$ is the dynamic range of the pixel-values typically chosen to be $12^{l}-$ 1 where $l$ is the number of bits per pixel.
@@
D.2 Distributional Metrics Validation
In this section, we explore the effectiveness of the proposed novel material distributional metrics (Sec. 4.2) by examining the nearest neighbor materials under certain distances.
In Fig. 12, we demonstrate the nearest reference BRDF to each synthetic sample under the distance
@@ D.2 Distributional Metrics Validation
Figure 12: Nearest reference under $d_{\text{log}}$ (Eq. 36): for each synthetic material in (a), we identify its nearest neighbor in the refernece set under $d_{\text{log}}$ presented at the same grid position in (b). The nearest neighbor metric is capable of matching most BRDFs in terms of reflective behaviors.
In Fig. 13, we further visualize the nearest neighbor information by plotting the heatmap of pairwise mean squared logarithmic distance defined as
@@
E Further Results
E.1 Unconditional Synthesis
For the unconditional synthesis experiment in Sec. 4.3, further results are presented in Figs. 14, 15, 16, 17 and 18 for NeuMaDiff (our method), PCA-A, PCA-N, VAE-A, and VAE-N, respectively
Figure 14: Synthesized materials by NeuMaDiff (our method).
@@ E.1 Unconditional Synthesis
F Additional Experiments
F.1 Statistical Analysis on MERL Dataset
We perform the statistical analysis on the MERL dataset [34]. We collect mean and maximum reflectance values in different color channels across all materials. Figure 19 demonstrates the statistical summary averaged across each material type while Fig. 20 details the statistics in each individual material grouped by type. In addition, to identify the classification boundaries between different material types, in Fig. 21, we apply K-means clustering (for background see Sec. A.5) on the maximum reflectance of red channel.
@@ F.1 Statistical Analysis on MERL Dataset
Figure 20: Full MERL BRDF [34] statistical analysis grouped by material type.
@@
F.2 Superresolution for Low-density BRDF
Experiment on the BRDF reconstruction for low-density input, where the position samples density are decreased $x$-time in all three dimension $\theta_{h},\theta_{d},\phi_{d}$ from $90\times 90\times 180$. We take $x=1,2,4,8,16$ as the scaling factor. Given each dimension $x$-time lower density, the number of position samples left are
$(1+\lfloor\frac{89}{x}\rfloor)\times(1+\lfloor\frac{89}{x}\rfloor)\times(1+\lfloor\frac{179}{x}\rfloor)$
The reconstruction baseline is to fetch BRDF values taking in position indices, which are rounded down to the nearest neighbor. On the other hand, NBRDF model is trained over low-density data. Under the same scene setting, I measure SSIM of rendered images between full-density ground truth and reconstructed BRDF (Tab. 2). Top results among the two methods are bold.

### v2

@@ B.2 Diffusion Model
During inference, the network enables sampling via an iterative process (song2022ddim), leveraging the factorization of the learned distribution as
@@
for $p({\bm{\mathbf{x}}}_{T}):=\mathcal{N}(0,I)$. For conditional sampling, we employ classifier-free guidance (CFG) (ho2022cfg) (please refer to Sec.˜D.2 for the algorithm).
B.3 Attention Mechanism
The attention mechanism (vaswani2017attention) allows models to focus on specific parts of input data, dynamically assigning different levels of “attention” or importance to different elements. The attention mechanism enables models to learn which parts of the input sequence are most relevant to predicting each output token. This process improves performance by allowing models to prioritize relevant information and ignore irrelevant details, especially in long sequences.
In modern applications, attention modules are widely used in natural language processing (NLP), computer vision, and beyond. The transformer architecture (vaswani2017attention), which relies heavily on the self-attention mechanism, has become foundational in NLP models, including BERT (devlin2018bert) and GPT (mann2020language). Attention allows these models to capture complex dependencies between words in a sentence, regardless of their distance from each other, leading to significant advancements in tasks like language translation, sentiment analysis, and image processing.
The attention mechanism computes a weighted combination of values based on the relevance of each value to a given query. In the context of self-attention (or scaled dot-product attention) in transformer models (vaswani2017attention), this process involves three main components: queries ($\bm{\mathbf{q}}$}), keys ($\bm{\mathbf{k}}$}), and values ($\bm{\mathbf{v}}$}). In summary, the attention mechanism dynamically focuses on relevant parts of the input by computing similarity scores between queries and keys, normalizing these scores, and using them to weight the values. This allows models to capture dependencies across elements in a sequence, making attention a powerful tool for handling long-range dependencies in data.
@@ B.3 Attention Mechanism
where $W_{\bm{\mathbf{q}}}$, $W_{\bm{\mathbf{k}}}$, and $W_{\bm{\mathbf{v}}}$ are learnable weight matrices. These projections represent the queries, keys, and values, respectively. The attention score between each query and key is then computed as the dot product ${\bm{\mathbf{q}}}\cdot{\bm{\mathbf{t}}}$. This results in a matrix of scores that represents the similarity between each element in the sequence. To stabilize gradients and prevent large values in the dot-product computation, the scores are scaled by the square root of the dimensionality of the queries/keys $\sqrt{d_{\bm{\mathbf{k}}}}$. The scaled scores are $\frac{{\bm{\mathbf{q}}}\cdot{\bm{\mathbf{t}}}}{\sqrt{d_{k}}}$.
@@
B.4 Variational Autoencoder
We adopt variational autoencoders (VAEs) (kingma2013auto) as one of the baseline models in Sec.˜4.3. VAEs are probabilistic generative models designed to capture the underlying probability distribution of a given dataset $\mathcal{D}$.
@@ B.4 Variational Autoencoder
$\mathcal{L}_{\text{ELBO}}({\bm{\mathbf{\psi}}},{\bm{\mathbf{\zeta}}};{\bm{\mathbf{x}}})$= $\mathbb{E}{q_{{\bm{\mathbf{\psi}}}}({\bm{\mathbf{z}}}|{\bm{\mathbf{x}}})}\left[\log p_{{\bm{\mathbf{\zeta}}}}({\bm{\mathbf{x}}}|{\bm{\mathbf{z}}})\right]$ (20)
@@
where $\mathcal{D}_{\text{KL}}$ represents the Kullback-Leibler divergence between the approximate posterior and the prior distribution (csiszar1975divergence). This probabilistic framework enables VAEs to effectively approximate the true data distribution, facilitating robust generative modeling of complex data.
B.5 K-Means Clustering
We adopt K-means clustering in constrained synthesis (Sec.˜4.5), where unconditional materials are filtered out by statistical information rather than a neural network. K-means clustering is a commonly used unsupervised method to partition $n$ data samples ${\bm{\mathbf{X}}}\in\mathbb{R}^{n\times p}$ into $k$ clusters $\mathcal{S}=\{S_{1},...,S_{k}\}$, minimizing the distance between each sample and its center
@@ B.5 K-Means Clustering
Appendix C Model Implementation Details
C.1 Neural Field
The dimensionality of the flattened weights for our neural field $f_{r}^{\bm{\mathbf{\xi}}}$ is $D_{\text{NF}}=675$.
C.2 Transformer Backbone in Hyperdiffusion
@@
Fig.˜10 illustrates the transformer architecture.
@@
C.3 PCA-Based Baselines
For PCA-based baselines introduced in the unconditional synthesis in Sec.˜4.3, the dimensionalities of the reduced spaces for PCA-A and PCA-N are 300 and 100, respectively.
C.4 VAE-Based Baselines
In this section, we detail the model hyperparameters for VAE-based baselines introduced in the unconditional synthesis in Sec.˜4.3.
For VAE-A, to reduce the input complexity, we downsample it from $90\times 90\times 180$ to $45\times 45\times 90$ and upsample the synthetic results using nearest neighbor algorithm. The hidden dimension is 256 and the latent space dimension is 300.
For VAE-N, the VAE architecture contains MLP-based encoder and decoder, each with 4 layers. The input dimension is $D_{\text{NF}}$, and the hidden and the latent space dimensions are both 300. The likelihood of the reconstructed data is measured by the mean squared error (MSE).
Appendix D Experiment Details
D.1 Training Details
In the neural field fitting, we set the batch size to 512, the number of epochs to 100, and the learning rate to $5\times 10^{-3}$.
When training the hyperdiffusion model, we set the total timestep $T$ to 100, the batch size to 512, the number of epochs to 700 for unconditional synthesis and an additional 200 for conditional synthesis, and the learning rate to adaptive from $5\times 10^{-4}$ to $5\times 10^{-6}$. The experiments are run on an NVIDIA GeForce RTX 4090 GPU.
D.2 Conditional Sampling with Classifier-Free Guidance
We employ classifier-free guidance (CFG) (ho2022cfg) for conditional sampling on the hyperdiffusion. We present the algorithm in Alg.˜1. Notice that we require $\omega>-1$ for conditional synthesis. Otherwise, the model downgrades to unconditional synthesis when $\omega=-1$.
@@ D.2 Conditional Sampling with Classifier-Free Guidance
${\bm{\mathbf{x}}}_{T}\sim\mathcal{N}(0,I_{D_{\text{NF}}})$ $\triangleright$ sample ${\bm{\mathbf{x}}}_{T}$ from prior
for $t=T$ to 11 do
@@
D.3 Full Material Type List
We include the full list of 48 material types used in the type-conditional synthesis in Sec.˜4.4: acrylic, alum-bronze, alumina-oxide, aluminium, aventurnine, brass, wood, chrome-steel, chrome, colonial-maple, color-changing-paint, delrin, diffuse-ball, fabric, felt, fruitwood, grease-covered-steel, hematite, ipswich-pine, jasper, latex, marble, metallic-paint, natural, neoprene-rubber, nickel, nylon, obsidian, oxidized-steel, paint, phenolic, pickled-oak, plastic, polyethylene, polyurethane-foam, pvc, rubber, silicon-nitrade, soft-plastic, special-walnut, specular-fabric, specular-phenolic, specular-plastic, stainless-steel, steel, teflon, tungsten-carbide, two-layer.
D.4 Full Set of Constrained Synthesis Rules
@@
    $\lvert\{({\bm{\mathbf{\omega}}}_{o},{\bm{\mathbf{\omega}}}_{i})\in\mathbb{R}^{6}\mid\left\|f_{r}({\bm{\mathbf{\omega}}}_{o},{\bm{\mathbf{\omega}}}_{i})\right\|_{\infty}>f_{\text{diffuse}}\}\rvert<e,$ (23)
    where $\left\|\cdot\right\|$ is the $\ell^{\infty}$-norm that selects the maximum component from the three color channels of reflectance values.
@@
  • Specular materials: Through our statistical analysis, we identify two specular thresholds $f_{\text{specular}}^{(1)}:=100$ and $f_{\text{specular}}^{(2)}:=600$. The materials can be classified as low-, mid-, or high-specular if the maximum reflectance value fall in the range $f_{\text{diffuse}},f_{\text{specular}}^{(1)}),[f_{\text{specular}}^{(1)},f_{\text{specular}}^{(2)})$, and $f_{\text{specular}}^{(2)},\infty)$, respectively.
@@
  • Mirror-like materials: the specular lobe in the polar diagram is narrow. In order to quantify this, we observe that under the Rusinkiewicz reparametrization (Sec.˜3.2), the evaluation at $f_{r}(\theta_{{\bm{\mathbf{h}}}}=0,\theta_{{\bm{\mathbf{d}}}}=0,\varphi_{{\bm{\mathbf{d}}}}=0)$ is the peak of the lobe. Incrementing $\theta_{\bm{\mathbf{h}}}$ decreases the BRDF. The width $w$ of the lobe can then be defined as the value of $\theta_{\bm{\mathbf{h}}}$ for which the reflectance value drops to half of the the peak $\frac{f_{r}(0,0,0)}{2}$. Empirically, we found that if the width $w<w_{\text{mirror}}:=0.349$, the material exhibits mirror-like behavior.
Appendix E Metric Details
E.1 Rendering-Based Metrics
@@
SSIM (SSIM04) is a perception-based metric that measures the perceptual similarity of the two images. The computation of SSIM is based on three comparison measurements between the two images: luminance ($l$), contrast ($c$), and structure ($s$) defined as
$l(I_{1},I_{2}):=\frac{2\mu_{1}\mu_{2}+c_{1}}{\mu_{1}^{2}+\mu_{2}^{2}+c_{1}};$ (25)
@@
are the variables to stabilize the division with weak denominator where $k_{1}$ and $k_{2}$ are coefficients defaulted to 0.01 and 0.03, respectively, and $L$ is the dynamic range of the pixel-values typically chosen to be $2^{l}-1$ where $l$ is the number of bits per pixel.
@@
E.2 Distributional Metrics Validation
In this section, we explore the effectiveness of the proposed novel material distributional metrics (Sec.˜4.2) by examining the nearest neighbor materials under certain distances.
In Fig.˜11, we demonstrate the nearest reference BRDF to each synthetic sample under the distance
@@ E.2 Distributional Metrics Validation
Figure 11: Nearest reference under $d_{\text{BRDF-L1-Log}}$ (Eq.˜35): for each synthetic material in (a), we identify its nearest neighbor in the reference set under $d_{\text{BRDF-L1-Log}}$ presented at the same grid position in (b). The nearest neighbor metric is capable of matching most BRDFs in terms of reflective behaviors.
In Fig.˜12, we further visualize the nearest neighbor information by plotting the heatmap of pairwise mean squared logarithmic distance defined as
@@
Appendix F Further Synthesis Results
For the unconditional synthesis experiment in Sec.˜4.3, further results are presented in Figs.˜13, 14, 15, 16 and 17 for M3ashy (our method), PCA-A, PCA-N, VAE-A, and VAE-N baselines, respectively. In addition, Fig.˜18 presents the complex geometries rendered with our synthesized materials. To support richer visual effects, our synthesized materials can also be rendered with normal maps (Fig.˜19), bump maps (Fig.˜20) and spatially varying configurations (Jakob2022DrJit) (Figs.˜21 and 22).
Figure 13: Synthesized materials by M3ashy (our method).
@@ Appendix F Further Synthesis Results
Figure 19: Our synthesized neural materials rendered with normal maps.
Figure 20: Our synthesized neural materials rendered with bump maps.
Figure 21: Our synthesized neural materials rendered in a spatially varying.
Appendix G Additional Experiments
G.1 Statistical Analysis on MERL Dataset
We perform the statistical analysis on the MERL dataset (Matusik2003datadriven). We collect mean and maximum reflectance values in different color channels across all materials. Figure˜23 demonstrates the statistical summary averaged across each material type while Fig.˜24 details the statistics in each individual material grouped by type. In addition, to identify the classification boundaries between different material types, in Fig.˜25, we apply K-means clustering (for background see Sec.˜B.5) on the maximum reflectance of red channel.
@@ G.1 Statistical Analysis on MERL Dataset
Figure 24: Full MERL BRDF (Matusik2003datadriven) statistical analysis grouped by material type.
@@
G.2 Superresolution from Low-Density BRDF
We present the experiment on the BRDF reconstruction for low-density input using NBRDF (sztrajman2021nbrdf), showing its capability of superresolution. In the experiment, when fitting the NBRDF to a MERL material (Matusik2003datadriven), we downsample the input $\theta_{h},\theta_{d},\phi_{d}$ by a factor of $x=1,2,4,8,16$ in all three angles. Since the original sampling density is $90\times 90\times 180$, after downsampling, the number of input samples is
$(1+\lfloor\frac{89}{x}\rfloor)\times(1+\lfloor\frac{89}{x}\rfloor)\times(1+\lfloor\frac{179}{x}\rfloor).$
For comparison we develop a baseline model, which just evaluates the BRDF according to the nearest neighbor. On the other hand, NBRDF model is trained over this downsampled data. Under the same scene setting, we compare SSIM of rendered images using the full-density groundtruth, the baseline, and the NBRDF model in Tab.˜4.